OpenStudy (unklerhaukus):

\[\large\lim_{x\to0^+}\frac{\sqrt x\sin(3x)+x^2+\arctan(5x)}{\sin\Big(\arctan\big(\sin(x)\big)\Big)}\]

Miracrown (miracrown):

have you tried plugging in 0 yet?

mathslover (mathslover):

That will not work actually ... Mira, don't u think it will lead to 0/0 form ?

OpenStudy (unklerhaukus):

\[\large\leadsto \tfrac00\]

Miracrown (miracrown):

yes it is 0/0 we plug in so that we know its of that form and so we can use lhospitals rule because it is 0/0

OpenStudy (unklerhaukus):

taking derivatives might get messy

Miracrown (miracrown):

yes but thats the only way to proceed

OpenStudy (dan815):

plit it into 3 term atleast who knows some of them might end up going to inf if u are lucky and u can stop

Miracrown (miracrown):

we have to use lhospitals rule here we just need to get the derivative of the top and the bottom that's what lhospitals rule is about we can use it for 0/0 or inf/inf for those cases

OpenStudy (unklerhaukus):

What derivatives do You get ?

mathslover (mathslover):

\(\lim_{x\rightarrow 0^+}\left(\cfrac{\cfrac{\sqrt{x} \sin (3x)}{\sqrt{1 - x \sin^2 (3x) }} + \cfrac{x^2}{\sqrt{1-x^4}} + \cfrac{\tan^{-1}(5x)}{\sqrt{1-\tan^{-1}(5x)^2}}}{\sin x}\right)\)

OpenStudy (unklerhaukus):

\[\leadsto\frac00\]

OpenStudy (dan815):

1 more lhopital gogo

mathslover (mathslover):

Yeah.. its not complete yet!

OpenStudy (dan815):

ganeshie do u like my bebe

mathslover (mathslover):

Dan, LH rule is not going to satisfy us. I will use something different here :-)

OpenStudy (dan815):

i was wondering why we don split that equation into 3 separate terms

OpenStudy (dan815):

and apply them separetly

OpenStudy (unklerhaukus):

do you want to see a simpler method?

OpenStudy (dan815):

yes!

mathslover (mathslover):

No..

OpenStudy (dan815):

gimme hint first

mathslover (mathslover):

I'm trying.. please.. Unkle, PM the hint to dan.

mathslover (mathslover):

Don't post it here :P I have already wasted 3 pages... I want to do more work ...

OpenStudy (dan815):

/haha math

mathslover (mathslover):

\[\large\lim_{x\to0^+}\frac{\sqrt x\sin(3x)+x^2+\arctan(5x)}{\sin\Big(\arctan\big(\sin(x)\big)\Big)}\] I copied the question again so that I don't have to scroll again and again. :P

mathslover (mathslover):

I really love that x^2 .. it looks the best! :P the others are nasty.

OpenStudy (mathmate):

What if we expand each bit by Taylor series and cancel the high order terms to get 0+0+5=5

OpenStudy (dan815):

aw math xD

OpenStudy (dan815):

mathmale got it

mathslover (mathslover):

That is what I tried... the other two are zero in my equation. The third expression should be 5.

OpenStudy (dan815):

the rest went to 0

OpenStudy (unklerhaukus):

@mathmate \[\large\lim_{x\to0^+} \frac{\sqrt x\sin(3x)+x^2+\arctan(5x)} {\sin\Big(\arctan\big(\sin(x)\big)\Big)}\] \begin{align*} &=\lim_{x\to0^+}\frac{\sqrt x\left[3x+O(x^3)\right]+x^2+\left[5x+O(x^3)\right]} {\sin\Big(\arctan\big(\left[x+O(x^3)\right]\big)\Big)} \\ &=\lim_{x\to0^+}\frac{3x^{3/2}+O(x^{5/2})+x^2+5x+O(x^3)} {\sin\Big(\left[x+O(x^3)\right]\Big)} \\ &=\lim_{x\to0^+}\frac{5x+3x^{3/2}+x^2+O(x^3)} {\left[x+O(x^3)\right]} \\ &=\lim_{x\to0^+}\frac{x(5+3x^{1/2}+x+O(x^2))} {x\big(1+O(x^2)\big)} \\ &=\lim_{x\to0^+}\frac{5+3x^{1/2}+x+O(x^2)} {1+O(x^2)} \\ &=\frac51 \\ &=5 \end{align*}

hartnn (hartnn):

\(\Large \sin (\arctan (\sin x)) \to \sin x \\ as \ x\to 0 \)

hartnn (hartnn):

then just separate the denominator

hartnn (hartnn):

and each ofthe 3 limits will be easy to solve

hartnn (hartnn):

***separate the numerator

hartnn (hartnn):

\(\Large \sin (\arctan (\sin x)) = \dfrac{\sin x}{\sqrt{1+\sin ^2 x}}\)

hartnn (hartnn):

\(\sqrt{1+\sin^2 x} \to 1 ~ as x\to 0\)

hartnn (hartnn):

sqrt x sin 3x /sin x -> 0 x^2/sin x -> 0 arctan 5x /sin x -> 5 as x-?0

mathslover (mathslover):

\(\color{blue}{\text{Originally Posted by}}\) @mathslover \(\lim_{x\rightarrow 0^+}\left(\cfrac{\cfrac{\sqrt{x} \sin (3x)}{\sqrt{1 - x \sin^2 (3x) }} + \cfrac{x^2}{\sqrt{1-x^4}} + \cfrac{\tan^{-1}(5x)}{\sqrt{1-\tan^{-1}(5x)^2}}}{\sin x}\right)\) \(\color{blue}{\text{End of Quote}}\) = 0 + 0 +5 Right!

OpenStudy (mathmate):

@UnkleRhaukus I always read your problems with awe and had no clue whatsoever. I am so glad even to think about participating this time.

OpenStudy (unklerhaukus):

Taylor series is powerful,

mathslover (mathslover):

I need to learn that *taylor series* .. it spoiled my work :/ Anyways, awesome question Unkle!!!

Miracrown (miracrown):

 @mathslover, any function that is regular, that means that all its derivative exist and are continuous in an interval can be expressed as a series of power